. Bill is killed in a car accident and his family is suing the other driver for the value of his lost wages. The man earned $60,000 per year and was expected to work for 4 more years (including the current year). You can treat his salary as being paid at the beginning of each year and the borrowing and lending rate of interest is 5% every year. a. If Bill’s salary is expected to remain the same and he will not die by some other causes before he retires, what is the PDV of his lost earnings? b. Suppose Bill’s salary is expected to grow by 5% per year and the probability that he dies from some other cause before year t is 2t (t=0 for the current year). What is the PDV of his lost earnings? c. The lawyer for the other driver wants to argue that the value of Bill’s lost earnings are lower than those calculated by Bill’s lawyer. Discuss two arguments the lawyer might give to justify his claim.
Answer #1

Answer for question 1

PDV=(60000*1.05/1.05)+(60000*1.05^2/1.05^2)+(60000/1.05^3)+(60000/1.05^4)=57142+54421+51830+49362=212755

Answer for 2

If he dies in year 1 probability of death is 2% in 1 year, 4% in 2nd year , 6 % in 3rd year , 8% in 4th year

He dies in 1st year or he dies in 2nd year or in 3rd or in4th year

Then PDV can be calculated as

PDV = 2% ((60000+(60000)+(60000)+(60000)) + 4%((60000)+(60000)+(60000))+6%((60000)+(60000))+8%(60000/)=2%*240000+4%180000+6%120000+8%60000=4800+7200+7200+4800=24000

As Growth rate equals to rate of return hence 60000 will remains the same notional amount as a salary.

Answer for c)

Driver lawyer can speak about his mortality rate he can advocate that his mortality chances are higher from other factors..

Rate of return considered by Bill s lawyer is less than the actual rate of return if he can argue that 5% is much less than real rate of return and growth rate as well